how do you write (5,7) and (-7,5) into a general form linear equation?
Answer by LinnW(1048)

First find the equation in slope intercept ( y = mx + b ) form
The first step is to find the slope m.
m = (y2-y1)/(x2-x1) where (x1,y1) = (5,7) and (x2,y2) = (-7,5)
m = (5 - 7)/(-7 -5)
m = (-2)/(-12)
m = 1/6
Now substitute 1/6 for m , 5 for x and 7 for y in y = mx +b
7 = (1/6)(5) + b
7 = 5/6 + b
add -5/6 to each side
6 1/6 = b
37/6 = b
Our equation is y = (1/6)x + 37/6
Now we need to change to general form.
Multiply each side by 6
6y = x + 37
add -x to each side.
-x + 6y = 37